OpenStudy (anonymous):

Felix exclaims that his quadratic with a discriminant of −1 has no real solutions. Felix then puts down his pencil and refuses to do any more work. Create an equation with a negative discriminant. Then explain to Felix, in calm and complete sentences, how to find the solutions, even though they are not real.

OpenStudy (anonymous):

-1

OpenStudy (mrnood):

Do you know what the discriminant is? If oyu write out the quadratic formula which part is the discriminant?

OpenStudy (anonymous):

b^2 - 4ac

OpenStudy (mrnood):

OK - and that appears in the "squareroot" sign It tells you tha ththe discriminant is -1 and we have a special name for sqrt(-1)... which is??

OpenStudy (anonymous):

I don't know?

OpenStudy (mrnood):

The question is about non real roots - also known as complex roots What is the basis of complex numbers?

OpenStudy (anonymous):

I don't know that either..

OpenStudy (mrnood):

OK - then I'm afraid you do not have the necessary background to answer this question. You seem to be on the subject of complex numbers - so maybe you need to refer back to you text or class notes on this subject....

OpenStudy (anonymous):

Would it be i?

OpenStudy (mrnood):

yes

OpenStudy (mrnood):

so \[\sqrt{-x} = \sqrt{-1 \times x} = \sqrt{-1} \times \sqrt{x}\]

OpenStudy (mrnood):

so write that with i in it